There are 2 copy machines. One can do a printing job in 75 minutes. If both machines print at the same time they can do the printing job in 50 minutes. How long will it take the other machine to do the printing job by itself.
Answer by ptaylor(1894)


Machine 1 can do 1/75 of the print job per minute
Machine 2 can do x of the print job per minute
Both together can do 1/50 of the print job per minute
So our equation to solve is:
1/75+x=1/50 multiply each term by 150
2+150x=3 subtract 2 from both sides
150x+2-2=3-2
150x=1 divide both sides by 150
x=1/150 of the print job per minute or
Machine #2 can do the print job by itself in 150 minutes---ans

Hope this helps------ptaylor